Problem:
Two angles are supplementary. Three times the smaller is five less than twice the larger. Find the number of degrees in each angle.
: Problem:
Two angles are supplementary. Three times the smaller is five less than twice the larger. Find the number of degrees in each angle.

Answer by josmiceli(2011)

Supplementary means the angles add up to 180degrees
Call the larger angle a
Call the smaller angle 180 - a
3*(180 - a) = 2a - 5
540 - 3a = 2a - 5
5a = 545
a = 109
180 - a = 71
The angles are 71 degrees and 109 degrees
check:
3*(180 - a) = 2a - 5
3*(180 - 109) = 2*109 - 5
3*71 = 218 - 5
213 = 213
OK